# PointDreamer: Zero-shot 3D Textured Mesh Reconstruction from Colored Point Cloud by 2D Inpainting

### v1

@@ B.1. Effect of Different $K$ Values (Number of Viewpoints for Projection and Inpainting)
6 At the centers of each face of a cube 20.8755 0.9273 0.0585 33.2870
Figure 14. Visual comparisons of our reconstructed meshes with different $K$ values (numbers of viewpoints for projection and inpainting). An insufficient number of views would lead to artifacts in invisible or occluded areas.
B.2. Visual Comparisons: Impact of Degraded Input Quality
B.2.1. Anti-Noise Ability Analysis
Figure 15 presents the visual comparisons of our method and baseline methods’ reconstructed meshes from noisy or clean input point clouds. Overall, we have the following observations, which are consistent with the quantitative results in our manuscript:
  1. (1)
  2. (2)
  3. (3)
Figure 15. Visual comparisons of our PointDreamer’s and baseline methods’ reconstructed meshes, with noisy or clean point clouds as input. Our PointDreamer shows a strong anti-noise ability by producing high-quality textures even with noisy input.
B.2.2. Sparsity Test
Figure 16. Visual comparisons of our PointDreamer’s reconstructed meshes with different numbers of points as input. There is a small performance drop when adopting sparser input, which sometimes can be hard to notice by human eyes.
B.3. Quantitative Results of Method Component Analysis
B.3.1. Geometry Extraction
We provide the quantitative results of replacing our geometry extraction module from POCO to SPR and Depth Inpainting in Table B.3.1. As can be seen, POCO, as a state-of-the-art deep-learning-based surface reconstruction approach, outperforms the other two methods by a significant margin, especially regarding FID. This is consistent with the visual comparisons (see Figure 9) in our manuscript where both SPR and Depth Inpainting suffer from noisy geometry and thus low-quality textures.
Table 7. Quantitative results of replacing our geometry extraction module from POCO to SPR and Depth Inpainting.
@@ B.3.1. Geometry Extraction
B.3.2. 2D Inpainting
We provide the quantitative results of replacing our 2D inpainting module from DDNM (Wang et al., 2022b) to nearest interpolation, linear interpolation, and DiffPIR (Zhu et al., 2023) in Table B.3.2. DiffPIR is another 2D-diffusion-based image restoration approach with inpainting ability. As can be seen, DDNM outperforms the other three methods thanks to its strong inpainting ability, which is consistent with the visual comparisons (see Figure 10) in our manuscript.
Table 8. Quantitative results of replacing our inpainting module from DDNM to nearest interpolation, linear interpolation, and DiffPIR.

### v2

@@ Appendix F Applications for large scenes
c—c—ccc ShapeNet Cat.   Method CD ↓ NC ↑ FS ↑
Chair   SPR 1.3094 0.9024 0.8710
NKSR 0.9079 0.8694 0.8188
POCO (TF & Ours) 0.4349 0.9267 0.9632
Car   SPR 0.8428 0.8847 0.8945
NKSR 0.7876 0.8201 0.8128
POCO (TF & Ours) 0.4407 0.8677 0.9462
Motirbike   SPR 4.2807 0.7647 0.6587
NKSR 0.8786 0.6871 0.7249
POCO (TF & Ours) 0.4423 0.7663 0.9434
Dataset   Method CD ↓ NC ↑ FS ↑
GSO   SPR 0.8565 0.9446 0.9265
NKSR 0.6149 0.9296 0.9006
POCO (TF & Ours) 0.4424 0.9474 0.9679
OmniObject3D   SPR 0.6777 0.9511 0.9407
NKSR 0.6148 0.9424 0.9003
POCO (TF & Ours) 0.3393 0.9669 0.9866
TABLE VI: Quantitative results of replacing our geometry extraction module from POCO to SPR and Depth Inpainting.
@@
TABLE VII: Quantitative results of replacing our inpainting module from DDNM to nearest interpolation, linear interpolation, and DiffPIR.
@@
Figure 9: Our reconstruction results on scenes.
Figure 10: More visual comparisons with baseline methods. Row 1-3: ShapeNetV2 dataset. Row4-6: GSO dataset. Row 7-9: OmniObject3d dataset.
Figure 11: Visual results with intermediate multi-view images.
Figure 12: Pipeline of extracting geometry by depth inpainting.
TABLE VIII: Quantitative results of using different $K$ values (numbers of viewpoints for projection and inpainting) on the motorbike category of ShapeNetCoreV2 dataset. More views contribute to a slightly higher reconstruction quality.
{NiceTabular}
6 At the centers of each face of a cube 20.8755 0.9273 0.0585 33.2870
Figure 13: Visual comparisons of our reconstructed meshes with different $K$ values (numbers of viewpoints for projection and inpainting). An insufficient number of views would lead to artifacts in invisible or occluded areas.
Figure 14: Visual comparisons of our PointDreamer’s and baseline methods’ reconstructed meshes, with noisy or clean point clouds as input. Our PointDreamer shows a strong anti-noise ability by producing high-quality textures even with noisy input.
Figure 15: Visual comparisons of our PointDreamer’s reconstructed meshes with different numbers of points as input. There is a small performance drop when adopting sparser input, which sometimes can be hard to notice by human eyes.
Figure 16: Visual comparisons of the Texture Field baseline trained with different losses. “2D per-pixel” denotes rendering the generated textured mesh to multi-view 2D images, and then calculating the MSE loss between the rendered and GT images. “3D per-point” denotes calculating MSE loss between the predicted and GT colors of sampled 3D points.